Given the first 4 central moments of a random variable x are 0,0.566,0.393 and 0.756 respectively, find the moment coefficient of skewness and kurtosis and interpret them.


1
Expert's answer
2021-07-12T13:22:53-0400

Solution:

First central moment"=E(X)=\mu=0"

Second central moment"=E[(X-\mu)^2]=E[(X-0)^2]=E[(X)^2]=0.566"

Third central moment"=E[(X-\mu)^3]=E[(X-0)^3]=E[(X)^3]=0.393"

Fourth central moment"=E[(X-\mu)^4]=E[(X-0)^4]=E[(X)^4]=0.756"

Now, "Var[X]=\sigma^2=" Second central moment"=0.566"

So, "\sigma=\sqrt{0.566}=0.752329"

Next, "Skewness(X)=\dfrac{E[(X-\mu)^3]}{\sigma^3}=\dfrac{0.393}{0.752329^3}=0.92293"

And, "Kurtosis(X)=\dfrac{E[(X-\mu)^4]}{\sigma^4}=\dfrac{0.756}{0.752329^4}=2.35988"

We know that for skewness is that if the number is greater than +1 or lower than –1, this is an indication of a substantially skewed distribution. Here, we get Skewness=0.92293 <+1, this shows there is no skewed distribution.

For kurtosis, we know that if the number is greater than +1, the distribution is too peaked.

Here, Kurtosis=2.35988>+1, this shows the distribution is too peaked.